OpenStudy (anonymous):

If v = 4.00 meters/second and makes an angle of 60° with the positive direction of the y-axis, what is the magnitude of vX?

OpenStudy (anonymous):

If I understood you correctly: the angle between the vector and the X-axis is 30˚. vX/v = cos(30) Vx= v*cos(30) = 4*cos(30) = 0.617 meters/second

OpenStudy (anonymous):

4 cos (30) = 3.46 m/s